A volleyball player serves a ball with his height of 1.5-meter height and

strikes the floor at point three meter horizontally.

a. Time of flight of the ball

b. Its horizontal velocity 


1
Expert's answer
2021-10-25T15:34:31-0400

Let v be the horizontal velocity. Initial velocity in the vertical direction is zero.

(a) So, time to reach the ground is given by,

"h = \frac{1}{2}gt^2"

"t = \sqrt{\frac{2h}{g}} = \sqrt{\frac{2 \times 1.5}{10}} = 0.55 s"


(b) Horizontal range is given by,

"x = vt"

"v = \frac{x}{t} = \frac{3}{0.55} = 5.45 m/s"